A uniform magnetic field strength 0.50T is directed at an angle of 60° to a plane of 100 turn,rectangular copper coil of length 0.04m and width 0.05m. The diameter of the copper wire which make up the coil is 1.00mm and resistivity is 1.7×10^-8 ohms metre.The magnetic field through the loop decreases to zero in 0.3seconds ,Determine
1)The initial flux through the coil
2)The average induced e.m.f in the coil as the field decreases
1
Expert's answer
2020-05-20T09:08:55-0400

1)


"\Phi=BA\cos{60}=0.5Bab\\=0.5(0.5)(0.05)(0.04)=0.0005\ Wb"

2)


"\epsilon=N\frac{\Phi}{t}=100\frac{0.0005}{0.3}=0.17\ V"